A manager has determined that a particular piece of equipment is likely to increase the cash flow of a business by $10,000 for each of the next five years. What is the present value of this equipment at an interest rate of 10%?

Please explain in detail
Answer #1

Cash flow of business rise by $10,000 each year

Rate of interest = 10% = 0.1

Present value of cash flow is determined by = Amount / (1 + rate of interest)Years after which payment is received

Assume first payment is received after the end of first year. Present value of first payment is 10,000 / (1 + 0.1)1 = 9,090,9

Present value of second payment is 10,000 / (1 + 0.1)2 = 8,264.46

Present value of third payment is 10,000 / (1 + 0.1)3 = 7,513.14

Present value of forth payment is 10,000 / (1 + 0.1)4 = 6,830.13

Present value of fifth payment is 10,000 / (1 + 0.1)5 = 6,209.21

Total present value of all cash flows = 6,209.21 + 6,830.13 + 7,513.14 + 8,264.46 + 9,090.9 = 37,907.84